if p(x)=x^3-3x^2+2x+5 and p(a)=p(b)=P(c)=0 then the value of (2-a)(2-b)(2-c)

Grade:9

2 Answers

ayush
7 years ago
P(x)=(x-a)(x-b)(x-c) as a,b,c are the roots thus putting x=2 in P(x) we get the value as 8-12+6+5=7 . Hope it`s correct ,please do answer.
ayush
7 years ago
I am sorry , it will be 8-12+4+5=5. Now hope it`s correct. ___________________Thank you.________________________